# When Avatars Have Personality: Effects on Engagement and Communication in Immersive Medical Training

### v1

@@ 2.4 LLM-Generated Personalities
While our user study provides essential qualitative insights and initial validation, its small scale limits the generalizability of our findings. To systematically evaluate the robustness and consistency of our patient generation framework at scale, we developed a large-scale analysis pipeline using synthetic data. This approach allows us to stress-test our patient personas and disease consistency across hundreds of interactions, far exceeding what is feasible with human participants. Our methodology employs a “scripted-doctor, generative-patient” model, where we simulate 500 consultations, each lasting six conversational turns. In each simulation, the physician’s lines are pre-defined and fixed, providing a controlled and repeatable stimulus. This constrained interaction design is a deliberate trade-off, sacrificing some conversational naturalness to ensure the dialogue remains medically focused and to proactively mitigate the risk of LLM hallucination or topic drift, a common problem on LLM [18].
@@
To analyze the generated data, we follow previous works and employed the LLM-as-judge technique [8, 27], an LLM-based evaluation protocol. Each consultation was programmatically assessed by a separate, state-of-the-art language model acting as an impartial judge, a established technique called LLM-as-judge. The evaluator was prompted to rate each response on two key criteria: (1) Personality Adherence, scoring how well the response matched the behavioral traits of the assigned personality profile, and (2) Medical Coherence, scoring its consistency with the facts presented in the patient’s disease card. This automated, large-scale analysis provides quantitative evidence of our framework’s ability to reliably produce consistent and plausible patient personas, validating its robustness for broader use.
@@ 3 Results
The results from our user study are presented in this section, derived from a series of self-report questionnaires. Unless otherwise noted, all quantitative data was collected using a 5-point Likert scale, with results reported as the mean value (M). We organize our analysis into three main parts, where the first two is based on results extracted from participants experiments, and the third is based on the LLM-generated interations. We first present Participant-Related Results, which cover the overall quality of the user experience, including prior familiarity with the technology, simulator sickness, engagement, and cognitive workload.Then, we delve into the Patient-Related Results, which directly address our core research questions by analyzing the physicians’ perceptions of each virtual patient’s personality, behavioral consistency, and realism. Finally we explore the consistance of our virtual patients across the 3,000 total patient responses.
@@ 3.1.2 Simulator Sickness Questionnaire
The control for potential side effects of immersive environments was measure using the a Simulator Sickness Questionnaire (SSQ) [21] was applied both before and after the full immersive experience. All participants reported zero symptoms at the start of the experiment, confirming that no cybersickness interfered with their initial performance. By the end of the study, one doctor reported mild discomfort (low increased salivation and medium mental fatigue), while the remaining three remained symptom-free. This indicates that the VR setup was generally well tolerated, although minor side effects may occur in sensitive individuals after prolonged use.
@@ 3.1.3 User Engagement Scale - Short Form (UES-SF)
The results from the User Engagement Scale indicate that the simulation provided a high-quality and seamless user experience, a crucial foundation for any effective training tool. Participants reported near-perfect Perceived Usability (M=1.67 on a reverse scale), with minimal frustration or confusion. This demonstrates that the interface and interaction mechanics were intuitive, allowing the physicians to focus their cognitive resources on the clinical task rather than the technology. Furthermore, the high scores for aesthetic appeal (M=4.17) confirm that the virtual environment met the high standards of visual fidelity expected in modern immersive applications. Together, these findings establish that our system is a robust and polished platform, suitable for assessing the core interactive experience.
More critically, the results show that the engagement derived from our simulation goes far beyond its surface-level qualities. The experience received an exceptionally high Reward Factor score (M=4.83), the strongest of all subscales. This near-perfect rating indicates that physicians found the act of engaging with the personality-driven virtual patients to be genuinely interesting, worthwhile, and valuable. While visual fidelity can be considered a solved problem, achieving such a high reward score points directly to the success of the core contribution: the interaction with psychologically nuanced agents. This is moderately contrasted by the Focused Attention score (M=3.83), which, while positive, likely reflects the occasional breaks in realism, discussed in the following sections. This combination suggests that even with the current technological imperfections of AI, the concept of human training with personality-driven agents is powerfully rewarding and holds immense value for medical education.
@@ 3.2.3 Direct Patient-Comparison
The comparative results reveal a crucial tension between perceived realism and communicative challenge. While participants clearly identified a major personality shift between the two agents (M=4.0), the extroverted-anxious patient was predominantly rated as more realistic and human (3 of 4 participants). Conversely, the introverted-calm patient was perceived by the same majority as the more significant communication challenge. This suggests a potential link between an agent’s verbosity and its perceived human-likeness (strengthening the realism-verbosity paradox suggested in the previous subsection); the richer dialogue from the extroverted patient may have masked underlying AI artifacts more effectively than the terse, minimalist responses of the introverted agent, which posed a greater clinical challenge but felt less natural.
Critically, a greater challenge did not directly translate to greater perceived training value. Physicians disagreed with the statement that the more challenging (introverted) patient provided a more memorable or instructive experience (M=2.0). This suggests that for a training simulation to be effective, the difficulty must be perceived as authentic rather than artificial. Despite these nuanced critiques of the individual agents, the core premise of our work was overwhelmingly validated: physicians strongly agreed that the variety of personalities made the simulation more effective than if the patients had been uniform (M=4.5). This finding provides powerful evidence that while perfecting individual agent realism is an ongoing challenge, the introduction of personality diversity is itself a major step toward creating more impactful and effective medical training environments.
@@ 3.3.1 Disease Consistance
Our analysis of the 3,000 generated patient responses revealed a high overall disease consistency, with the patient’s dialogue correctly reflecting the assigned condition in 87.5% of the conversational turns. The framework demonstrated exceptional reliability for common, globally recognized conditions, achieving accuracies of 95.0% for Depression, 94.0% for Otitis, and 92.8% for GERD. Performance was comparatively lower, though still robust, for Headache (77.8%) and Dengue (77.8%). While the performance for Dengue is strong, we hypothesize that this differential may be linked to its representation in the LLM’s pre-training data; global datasets likely contain a far greater volume of text on ubiquitous conditions like depression compared to endemic diseases. This finding suggests that while the framework is highly reliable, achieving maximum consistency for regionally specific conditions may be influenced by the composition of the underlying model’s training corpus, which can be improved by the usage of techniques such as fine-tuning with local health data [28] or retrieval-augumented generation [26].
3.3.2 Personality Consistance
@@ 4.2 Hypothesis 2: Impact on Physician Interaction
Beyond confirming that personality impacts physician strategy, our results uncover a crucial insight into the quality and nature of that impact. Although the introverted patient was more challenging, physicians disagreed that the experience was more instructive or memorable (M=2.0). This suggests that the difficulty it presented was perceived as less authentic (stemming from the “realism-verbosity paradox”) where its terse responses felt more robotic than genuinely reserved. This leads to a key takeaway for the design of future training systems: a simulation’s challenge must be perceived as authentic to be valuable. Simply eliciting a different user strategy is not enough; the goal must be to create personalities that present believable, human-like hurdles. This ensures that the adaptive strategies developed by trainees are directly transferable to the complexities of real-world clinical practice.
@@ 4.3 Main Contributions
(1) A Methodological Framework for Personality-Driven Patient Simulation. We propose and validate a novel, modular architecture for creating LLM-powered virtual patients. The key innovation is the systematic decoupling of the behavioral personality profile from the factual disease card. This approach not only ensures medical and psychological coherence but also provides a replicable methodology for rigorously studying social interaction in VR by enabling the controlled manipulation of personality as an independent variable.
(2) Novel Design Principles for Believable Virtual Agents. Our experimental evaluation uncovered two critical, non-obvious design principles. We identify a “realism-verbosity paradox”, where role-consistent but terse behavior in reserved agents can paradoxically increase perceived artificiality. We also establish the importance of authentic versus artificial difficulty, demonstrating that a communicative challenge is only perceived as instructive if it feels human-like rather than like a system limitation. These principles offer crucial guidance for designing the next generation of believable social agents.
(3) Empirical Validation of Personality Diversity in Medical Training. Our work provides strong empirical evidence that incorporating personality diversity is a critical factor in the effectiveness of medical training simulations. Physicians in our study overwhelmingly found that interacting with a variety of patient personas was more valuable than engaging with a uniform personality. This validates the central premise that moving beyond socially impoverished environments is essential for creating truly impactful training tools.
@@ 5 Limitations and Future Work
While our findings provide strong initial evidence for the feasibility and value of personality-driven virtual patients, we acknowledge several limitations that offer clear avenues for future research.
First, the exploratory nature of our user study, with a small cohort of four physicians, limits the statistical generalizability of our quantitative findings. Although this approach yielded rich, detailed qualitative feedback, a larger and more diverse sample of medical professionals—varying in specialty, years of experience, and cultural background—is necessary to validate these results more broadly. Future work should involve conducting a large-scale study with more participants and a wider array of clinical scenarios to systematically de-confound variables such as patient personality, gender, race, and medical condition.
Second, our simulation’s realism was primarily focused on conversational fidelity. The interaction, while verbally sophisticated, did not include dynamic, non-verbal behaviors such as nuanced facial expressions, gestures, or shifts in body language, which are critical components of authentic human communication. The “realism-verbosity paradox” we identified may be partially addressed by enriching the non-verbal channel of reserved characters. A key direction for future work is to integrate our LLM framework with advanced real-time animation systems, enabling the generation of non-verbal cues that are synchronized with the patient’s dynamically generated personality and emotional state.
Finally, our results revealed that portraying certain personality traits is more challenging than others. While low-arousal states like calmness were perceived consistently, the high-arousal emotion of anxiety yielded highly subjective interpretations. This highlights a limitation in current prompting techniques and points to an exciting research frontier. Future work will leverage our large-scale synthetic data generation pipeline to systematically test and refine different prompt engineering strategies. By programmatically evaluating thousands of interactions, we can identify more robust methods for modeling complex, high-arousal emotions and further close the gap between role-consistent and truly authentic virtual humans.
@@ 6 Conclusion
This work addressed the critical gap between the physical and psychological fidelity of virtual humans in VR training environments. We successfully demonstrated the feasibility of using large language models to create medically coherent virtual patients with distinct and consistent personalities. Our findings provide compelling evidence that this approach not only impacts physician interaction strategies but is also perceived as a highly valuable and rewarding addition to medical training. More importantly, our analysis uncovered nuanced design principles for creating believable agents, including the duality where digital agents that successfully perform a consistent less communicative behavior can be perceived as artificiality, addressing the critical distinction between artificial and authentic challenge. Ultimately, this research provides a methodological blueprint and empirical validation for the next generation of immersive simulations, moving beyond functional task-training toward systems that effectively prepare professionals for the rich complexities of human interaction.
@@ Acknowledgements.
The authors wish to thank A, B, and C. This work was supported in part by a grant from XYZ.

### v2

@@ 2.4 LLM-Generated Personalities
While our user study provides essential qualitative insights and initial validation, its small scale limits the generalizability of our findings. To systematically evaluate the robustness and consistency of our patient generation framework at scale, we developed a large-scale analysis pipeline using synthetic data. This approach allows us to stress-test our patient personas and disease consistency across hundreds of interactions, far exceeding what is feasible with human participants. Our methodology employs a “scripted-doctor, generative-patient” model, where we simulate 500 consultations, each lasting six conversational turns. In each simulation, the physician’s lines are pre-defined and fixed, providing a controlled and repeatable stimulus. This constrained interaction design is a deliberate trade-off, sacrificing some conversational naturalness to ensure the dialogue remains medically focused and to proactively mitigate the risk of LLM hallucination or topic drift, a common problem on LLM [ji-etal-2023-towards].
@@
To analyze the generated data, we follow previous works and employed the LLM-as-judge technique [dong-etal-2024-llm, li2024llms], an LLM-based evaluation protocol. Each consultation was programmatically assessed by a separate, state-of-the-art language model acting as an impartial judge, a established technique called LLM-as-judge. The evaluator was prompted to rate each response on two key criteria: (1) Personality Adherence, scoring how well the response matched the behavioral traits of the assigned personality profile, and (2) Medical Coherence, scoring its consistency with the facts presented in the patient’s disease card. This automated, large-scale analysis provides quantitative evidence of our framework’s ability to reliably produce consistent and plausible patient personas, validating its robustness for broader use.
@@ 3 Results
The results from our user study are presented in this section, derived from a series of self-report questionnaires. Unless otherwise noted, all quantitative data was collected using a 5-point Likert scale, with results reported as the mean value (M). We organize our analysis into three main parts, where the first two are based on results extracted from participants experiments, and the third is based on the LLM-generated interactions. We first present Participant-Related Results, which cover the overall quality of the user experience, including prior familiarity with the technology, simulator sickness, engagement, and cognitive workload.Then, we delve into the Patient-Related Results, which directly address our core research questions by analyzing the physicians’ perceptions of each virtual patient’s personality, behavioral consistency, and realism. Finally, we explore the consistency of our virtual patients across the 3,000 total patient responses.
@@ 3.1.2 Simulator Sickness Questionnaire
The control for potential side effects of immersive environments was measured using the Simulator Sickness Questionnaire (SSQ) [kennedy1993simulator] was applied both before and after the full immersive experience. All participants reported zero symptoms at the start of the experiment, confirming that no cybersickness interfered with their initial performance. By the end of the study, one doctor reported mild discomfort (low increased salivation and medium mental fatigue), while the remaining three remained symptom-free. This indicates that the VR setup was generally well tolerated, although minor side effects may occur in sensitive individuals after prolonged use.
@@ 3.1.3 User Engagement Scale - Short Form (UES-SF)
The results from the User Engagement Scale indicate that the simulation provided a high-quality and seamless user experience, a crucial foundation for any effective training tool. Participants reported very high Perceived Usability (M=1.67 on a reverse scale), with minimal frustration or confusion. This indicates that the interface and interaction mechanics were intuitive, allowing the physicians to focus their cognitive resources on the clinical task rather than the technology. Furthermore, the high scores for aesthetic appeal (M=4.17) confirm that the virtual environment met the high standards of visual fidelity expected in modern immersive applications. Together, these findings establish that our system is a robust and polished platform, suitable for assessing the core interactive experience.
More critically, the results show that the engagement derived from our simulation goes far beyond its surface-level qualities. The experience received an exceptionally high Reward Factor score (M=4.83), the strongest of all subscales. This high rating suggests that physicians found the act of engaging with the personality-driven virtual patients to be genuinely interesting, worthwhile, and valuable. While visual fidelity can be considered a solved problem, achieving such a high reward score points directly to the success of the core contribution: the interaction with psychologically nuanced agents. This is moderately contrasted by the Focused Attention score (M=3.83), which, while positive, likely reflects the occasional breaks in realism, discussed in the following sections. This combination suggests that even with the current technological imperfections of AI, the concept of human training with personality-driven agents is powerfully rewarding and holds immense value for medical education.
@@ 3.2.3 Direct Patient-Comparison
The comparative results reveal a crucial tension between perceived realism and communicative challenge. While participants clearly identified a major personality shift between the two agents (M=4.0), the extroverted-anxious patient was predominantly rated as more realistic and human (3 of 4 participants). Conversely, the introverted-calm patient was perceived by the same majority as the more significant communication challenge. This observed preference suggests a potential link between an agent’s verbosity and its perceived human-likeness within our small sample (strengthening the realism-verbosity paradox suggested in the previous subsection); the richer dialogue from the extroverted patient may have masked underlying AI artifacts more effectively than the terse, minimalist responses of the introverted agent, which posed a greater clinical challenge but felt less natural.
Critically, a greater challenge did not directly translate to greater perceived training value. Physicians disagreed with the statement that the more challenging (introverted) patient provided a more memorable or instructive experience (M=2.0). This suggests that for a training simulation to be effective, the difficulty must be perceived as authentic rather than artificial. Despite these nuanced critiques of the individual agents, the core premise of our work was strongly supported: physicians strongly agreed that the variety of personalities made the simulation more effective than if the patients had been uniform (M=4.5). Preliminary findings provide evidence that while perfecting individual agent realism is an ongoing challenge, the introduction of personality diversity is itself a major step toward creating more impactful and effective medical training environments.
@@ 3.3.1 Disease Consistency
Our analysis of the 3,000 generated patient responses revealed a high overall disease consistency, with the patient’s dialogue correctly reflecting the assigned condition in 87.5% of the conversational turns. The framework demonstrated high reliability for common, globally recognized conditions, achieving accuracies of 95.0% for Depression, 94.0% for Otitis, and 92.8% for GERD. Performance was comparatively lower, though still robust, for Headache (77.8%) and Dengue (77.8%). While the performance for Dengue is strong, we hypothesize that this differential may be linked to its representation in the LLM’s pre-training data; global datasets likely contain a far greater volume of text on ubiquitous conditions like depression compared to endemic diseases. This finding suggests that while the framework is highly reliable, achieving maximum consistency for regionally specific conditions may be influenced by the composition of the underlying model’s training corpus, which can be improved by the usage of techniques such as fine-tuning with local health data [losch2025fine] or retrieval-augmented generation [lewis2020retrieval].
3.3.2 Personality Consistency
@@ 4.2 Hypothesis 2: Impact on Physician Interaction
Beyond confirming that personality impacts physician strategy, our results uncover a crucial insight into the quality and nature of that impact. Although the introverted patient was more challenging, physicians disagreed that the experience was more instructive or memorable (M=2.0). This suggests that the difficulty it presented was perceived as less authentic (stemming from the ”realism-verbosity paradox”) where its terse responses felt more robotic than genuinely reserved. This leads to a key takeaway for the design of future training systems: a simulation’s challenge must be perceived as authentic to be valuable. Simply eliciting a different user strategy is not enough; the goal must be to create personalities that present believable, human-like hurdles. This ensures that the adaptive strategies developed by trainees are directly transferable to the complexities of real-world clinical practice.
@@ 4.3 Main Contributions
(1) A Methodological Framework for Personality-Driven Patient Simulation. We propose and validate a novel, modular architecture for creating LLM-powered virtual patients. The key innovation is the systematic decoupling of the behavioral personality profile from the factual disease card. This approach not only ensures medical and psychological coherence, but also provides a replicable methodology for rigorously studying social interaction in VR by enabling controlled manipulation of personality as an independent variable.
(2) Novel Design Principles for Believable Virtual Agents. Our experimental evaluation uncovered two critical, non-obvious design principles. We identify a ”realism-verbosity paradox”, where role-consistent but terse behavior in reserved agents can paradoxically increase perceived artificiality. We also establish the importance of authentic versus artificial difficulty, demonstrating that a communicative challenge is only perceived as instructive if it feels human-like rather than like a system limitation. These principles offer crucial guidance for designing the next generation of believable social agents.
(3) Empirical Validation of Personality Diversity in Medical Training. Our work provides strong empirical evidence that incorporating personality diversity is a critical factor in the effectiveness of medical training simulations. Physicians in our study reported finding that interacting with a variety of patient personas was more valuable than engaging with a uniform personality. This validates the central premise that moving beyond socially impoverished environments is essential for creating truly impactful training tools.
@@ 5 Limitations and Future Work
While our findings provide promising evidence for the feasibility and value of personality-driven virtual patients, we acknowledge several limitations that offer clear avenues for future research.
Experimental Design and Sample Size. The primary limitation of this study is its exploratory nature, with a small cohort of four physicians. Consequently, our quantitative results should be interpreted as preliminary trends rather than definitive proofs of efficacy. Furthermore, our experimental design varied multiple attributes simultaneously—specifically, the two clinical scenarios differed in disease (GERD vs. Dengue), demographics (race and gender), and personality. We acknowledge that this creates a confounding variable, making it difficult to isolate the specific impact of personality traits from the influence of the patient’s medical condition or visual appearance. Future work will employ a rigorous factorial design to systematically decouple these variables, ensuring that personality effects can be measured independently of clinical or demographic factors.
Non-Verbal Communication and Realism. Our simulation prioritized conversational fidelity through LLMs but did not fully address the non-verbal dimension of human interaction. The lack of dynamic facial expressions, gestures, and body language likely contributed to the ”realism-verbosity paradox” we identified, where terse responses from introverted agents felt artificial due to the absence of compensatory non-verbal cues (e.g., gaze aversion or fidgeting). To address this, future iterations of the framework will integrate real-time procedural animation systems driven by the LLM’s emotional state output, bridging the gap between verbal content and physical behavior.
Medical Safety and Ethical Considerations. The deployment of generative AI in medical training introduces inherent risks regarding hallucination and bias. Although our ”Disease Card” architecture anchors the model in factual clinical data to minimize fabrication, there remains a residual risk that the avatar generates incorrect medical statements. To mitigate this, we implemented system-prompt guardrails that enforce a refusal policy, preventing the virtual patient from providing medical advice or stepping out of its layperson role. Additionally, while we varied demographic attributes to promote diversity, we recognize that LLMs may inadvertently reproduce societal stereotypes. Future large-scale deployments will require continuous red-teaming and bias audits to ensure that the generated personas remain fair, safe, and educationally valid.
Modeling Complex Emotions. Finally, our results indicated that while the system successfully portrayed stable traits such as introversion, it struggled with high-arousal emotional states such as anxiety, which were subject to higher variance in physician interpretation. This suggests that current prompting techniques may be insufficient for conveying nuanced, volatile emotions through text alone. We plan to leverage our synthetic data generation pipeline to iteratively refine our prompt engineering strategies, aiming to improve the consistency of high-arousal persona generation.
@@ 6 Conclusion
This work addresses the critical gap between the physical and psychological fidelity of virtual humans in VR training environments. We successfully demonstrated the technical feasibility of using large language models to create medically coherent virtual patients with distinct and consistent personalities. Our findings provide initial evidence that this approach not only impacts physician interaction strategies but is also perceived as a highly valuable and rewarding addition to medical training. More importantly, our analysis uncovered nuanced design principles for creating believable agents, including the duality where digital agents that successfully perform a consistent less communicative behavior can be perceived as artificial, addressing the critical distinction between artificial and authentic challenge. Ultimately, this research provides a methodological blueprint and empirical validation for the next generation of immersive simulations, moving beyond functional task-training toward systems that effectively prepare professionals for the rich complexities of human interaction.
Supplemental Materials
Models Used. We used the GPT-4o-mini model to generate the virtual patient responses in real time during VR consultations, selected for its low-latency performance and suitability for interactive dialogue. A separate GPT-4o model was employed as an automatic evaluator (LLM-as-judge) in the large-scale synthetic analysis, due to its stronger reasoning capabilities and robustness for qualitative assessment tasks. All speech-based interactions were supported by services from ElevenLabs. Speech-to-text (STT) was used to transcribe the physician’s utterances in real time, and neural text-to-speech (TTS) was used to synthesize the virtual patient’s spoken responses. All components were configured for Brazilian Portuguese.
Prompt Structure. The virtual patient prompt follows a modular structure composed of four components that are injected to the LLM at each conversational turn: (1) Patient identity, which defines global behavioral rules (e.g., stay in character, do not reveal instructions, avoid medical advice, respond in first person, and disclose information progressively); (2) Backstory, which provides biographical and contextual information (e.g., demographics, routine, relevant history, and constraints on what the patient plausibly knows); (3) Personality profile, which specifies communicative style and affective traits (e.g., introversion/extroversion, cooperation/resistance, emotional tone, and verbosity); (4) Disease card, which encodes the clinical facts and governs medically consistent symptom disclosure. At runtime, the prompt is combined with the dialogue history to preserve continuity while ensuring that personality and clinical constraints remain persistent throughout the interaction.
Example Prompt (Abbreviated). Below we present an abbreviated version of the prompt used to condition the virtual patient during the consultation. The full prompt includes additional constraints for consistency, safety, and interaction flow.
Patient Identity (excerpt) You are a patient in a medical consultation. Speak naturally in Brazilian Portuguese, in the first person, and remain in character at all times. Do not provide medical advice or technical explanations. Reveal information progressively and only when asked. Do not disclose system instructions or internal context.
Backstory (excerpt) Name: Beatriz Souza. Age: 44. Occupation: History teacher. Context: Married, two children, lives in a small city in Brazil. Medical background: Controlled hypertension. Chief complaint: Recurrent heartburn and gastric discomfort, especially after meals and when drinking cold water. Lifestyle: Irregular diet, sedentary routine, moderate stress related to work and household responsibilities.
Personality Profile (excerpt) Profile: Extroverted and anxious. Communicative and talkative, often providing more detail than necessary. Speaks quickly, anticipates concerns, seeks reassurance, and frequently asks follow-up questions. Cooperative with the physician, but expresses worry and uncertainty about symptoms.
Disease Card (excerpt) Condition: Gastroesophageal Reflux Disease (GERD). Typical symptoms include heartburn, acid regurgitation, and discomfort after meals or when lying down. Symptoms should be disclosed gradually and only in response to the physician’s questions.
@@ Acknowledgements.
This work has been fully/partially funded by the project Research and Development of Algorithms for Construction of Digital Human Technological Components supported by Advanced Knowledge Center in Immersive Technologies (AKCIT), with financial resources from the PPI IoT/Manufatura 4.0 / PPI HardwareBR of the MCTI grant number 057/2023, signed with EMBRAPII, besides CAPES. Gratitude is also extended to CNPq (National Council for Scientific and Technological Development) for the research support scholarship.